OpenStudy (anonymous):

. Balance these equations: (a) ____Fe(s) + ____O2(g)h____ Fe2O3(s) (b) ____H2(g) + ____N2(g)h____NH3(g) (c) ____Cl2(g) + ____KBr(aq)h____Br2(l ) +____KCl (aq) (d) ____CH4(g) + ____O2(g)h____CO2(g) +____H2O(l ) @diolle

OpenStudy (anonymous):

what's the formula in finding the amu?

OpenStudy (anonymous):

????

OpenStudy (anonymous):

can you help me answer this? The element nitrogen has an atomic weight of 14.0 and consists of two stable isotopes nitrogen-14 and nitrogen-15. The isotope nitrogen-14 has an atomic mass of 14.0 amu and a percent natural abundance of 99.6 %. The isotope nitrogen-15 has a percent natural abundance of 0.370 %. What is the atomic mass of nitrogen-15?________amu

OpenStudy (anonymous):

?

OpenStudy (anonymous):

14.0

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

oh it's 15

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

just solved it.

OpenStudy (anonymous):

i just solved what?

OpenStudy (anonymous):

\[[14.0 - 0.996 14.0]/3.70e=15.0 amu\] is that right?

OpenStudy (anonymous):

ok but hoe dp i balance these?

OpenStudy (anonymous):

sorry can't help you with that. I don't know that one

OpenStudy (anonymous):

yes that is right

OpenStudy (anonymous):

can anyone balance these equations i don't get it?

OpenStudy (anonymous):

Alright, so you just have to balance these so that they are equal on both sides: a) \[4Fe + 3O _{2} \rightarrow 2Fe _{2}O_{3}\] b) 3H2 + N2 = 2NH3 c) Cl2 + 2KBr = Br2 + 2KCl d) CH4 + 2O2 = CO2 + 2H20

OpenStudy (anonymous):

Like in a), all you have to do is find a common multiple between 3 and 2 (6) to balance your oxygen on both sides of the equation. So if you multiply Fe2O3 by 2, you will have 2FeO3 (2 Fe, and 6 O) Likewise, if you multiply O2 by a coefficent of 3, you will have 6 O on that side, thus balancing the oxygen. Then, simply balance the remaining Fe. There are 4 Fe (2 x 2) on the products side, so you must have a coefficient of 4 on the reactants side. Tip: It may be easier to balance your compounds first, especially ones containing O, and then tackle the elements and ions. Hope this helps!